An aerobatic aircraft is performing circular loops at the speed of sound, covering spectacularly 340 meters every single second in a flight exhibition. All participants have been warned that they keep their acceleration below 8 times that of the gravitational acceleration to avoid getting dizzy. What is the diameter of the smallest loop that these participants can safely do? Here you may use 10 m/s/s as the gravitational acceleration.


1
Expert's answer
2021-05-28T07:17:47-0400

Speed of aircraft in circular loop, "v=340\space m/s"

Maximum acceleration permissible "= 8g=80\space m/s^2"

Let the radius of the smallest loop "=R"

Diameter, "D=2R"

"R=\dfrac{D}{2}"

Centripetal acceleration, "a_c=\dfrac{v^2}{R}"

"R=\dfrac{v^2}{a_c}"

"D=\dfrac{2v^2}{a_c}=\dfrac{2\times(340)^2}{80}"

"D=2890\space m"